OpenStudy (anonymous):

Is the simplication of 3^8/3^3, 1/15?

OpenStudy (rea201):

3^8/3^3=3^8-3=3^5

OpenStudy (anonymous):



OpenStudy (jhannybean):

\[(\frac{ 3^8 }{ 3^3 }) = 3^{8-3} = 3^5 =243\]

OpenStudy (anonymous):

Oh, Thank you

OpenStudy (jhannybean):

When you have two exponents with the same base, and are dividing them, the powers will subtract.

OpenStudy (anonymous):

But doesn't 3 and 3 cancel out?

OpenStudy (jhannybean):

No you cannot cancel anything out until you have simplified them to the max

OpenStudy (anonymous):

Oh. Thank you

OpenStudy (jhannybean):

no problem :)